Grade 12th passModern Physics

A pair of eye glass frames is made of epoxy plastic at room temperature20.0°c, the frames have circular lens holes 2.20cm in radius. To what temperature must the frames be heated if lens 2. 21cm in radius are to be inserted in them? The average coefficient of linear expansion for epoxy is 1.30x10-4°c-1

Jessica oghre
8 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Recall that, by linear expansion: 
∆L = aL₁∆T. 

Solving this for ∆T gives: 
∆T = ∆L/(aL₁) 

Since the want the lenses to be 2.21 cm and they are initially 2.20 cm, we have: 
∆L = 2.21 - 2.20 = 0.01 cm. 

Plugging this into our equation gives: 
∆T = ∆L/(aL₁) 
= (0.01 cm)/[(1.30 x 10^-4° C^-1)(2.20 cm)] 
= 34.9°C. 

Therefore, the temperature desired is 20.0 + 34.9 = 54.9°C. 

I hope this helps!